please help me find the x intercepts of a parabola
Y=x^2-9x+18

Answer by scott8148(5897)

x-intercepts are when y=0 ___ also called roots

0 = x^2 - 9x + 18

factoring ___ 0 = (x - 6)(x - 3)

x - 6 = 0 ___ x = 6

x - 3 = 0 ___ x = 3